OpenStudy (anonymous):

Help with radicals please! 4

OpenStudy (alexandervonhumboldt2):

ok

OpenStudy (anonymous):

\[\sqrt{243}-\sqrt{75}=n\]

OpenStudy (anonymous):

How do I subtract them?

OpenStudy (anonymous):

you cannot

OpenStudy (alexandervonhumboldt2):

ok let me think hw to say it i know it

OpenStudy (anonymous):

Well then what is the value of n

OpenStudy (alexandervonhumboldt2):

9 sqrt3 - 5 sqrt3=4 sqrt3

OpenStudy (alexandervonhumboldt2):

you need to simplify them first then just

OpenStudy (anonymous):

Thanks